# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

### v2

@@ Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions. Tuning large arrays of QD qubits depends on efficient choices of automated protocols. Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization. For forming the current channels, BATIS follows a non-standard approach that requires a single measurement regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring only minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, it will no longer be possible to tune up many dots simultaneously using human-based techniques. The field of autotuning has arisen to address this challenge.
@@
Modern semiconductor QD qubit devices have two critical features that enhance their performance, while simultaneously making autotuning challenging and essential. First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping. All charge carriers—electrons or holes—are induced by gate voltages. No current can flow until gate voltages are applied and carriers are accumulated, yet turn-on voltages are unknown a priori. Furthermore, because of device drift and hysteresis, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some can host an array of QD qubits, while others provide a home for charge sensors. These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable. However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.. Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers, these channels must interact to enable the formation of QDs separated by tunnel barriers and proximal charge sensors.
Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including (critically) where current channels should be continuous and which gates should divide one channel from another. The outputs of BATIS are the gate voltages that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known. Importantly, in this work, the sample is kept at around —about 10 times warmer than other typical devices [10, 11]—which is highly desirable for high-temperature operation [12]. However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [13]. We show that BATIS works well at these high temperatures, which is important both for scalability and automated testing, e.g., using a cryogenic wafer prober [14, 15].
@@
Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases. For example, setting the device topology in a modulation-doped QD device with predefined current channels has been demonstrated in Refs. [16, 17, 18]. Autotuning has similarly been implemented in an undoped SiMOS QD to achieve single-electron occupation [19]. For nanowires, which by design have a single current channel, automated tuning into fully formed QDs has been demonstrated, provided device-specific details were preconfigured [20].
@@
Additionally, for undoped, 1D accumulation-mode QD devices, previous work has enabled progress in tuning a device topology [22, 23] and calibrating the charge state [16, 24]. Methods for automatically detecting unintentional, spurious QDs have also been proposed [25]. More recently, virtual-gate control of the electrostatic potential landscape confining holes within a 2D QD array has been demonstrated [26]. However, autonomous bootstrapping for accumulation-mode QD devices has never been attempted. Rather, most proposed tuning algorithms assumed that initial steps, such as device characterization and channel formation, have already been accomplished [27].
Our approach for BATIS differs from the existing bootstrapping algorithms, which often use efficiency as a figure of merit. By focusing on speed, such algorithms tend to be specialized to a particular device, allowing little room for benchmarking on different devices or in different experimental setups. Even if such algorithms perform well in one setting, they require a significant redesign to work in other settings. In contrast, we have specifically designed BATIS to be platform-agnostic, so that it may be applied to a variety of QD devices. While the procedural flow of BATIS is well-defined, the executable flow is automatically adjusted to suit the device architecture and external control lines, both of which are specified in the configuration file. In other words, rather than following a fixed sequence of measurements, BATIS—in real-time—chooses which steps to take and defines the expected outcomes depending on the device used. As the tuning progresses and the algorithm learns the device’s characteristics, BATIS applies its knowledge to automatically customize further measurements.
Finally, we note that existing autotuning schemes focus mainly on the more advanced stages of a tune-up. These include coarse tuning, charge-state tuning, virtualization, and fine-tuning, which are sufficient for QDs formed in a preexisting channel, or when the initial bootstrapping steps are performed by hand [27]. In modern chips, hand-tuning will soon be precluded as QD processors scale up. BATIS is designed to autonomously adjust to the QD layout and perform the initial tune-up stages, including checking for fabrication defects and characterizing local and global current flow, all of which are essential for large-scale qubit implementations. It thus addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
II BATIS Design
@@
Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming QDs (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [28, 29].
@@
The screening gates (yellow) are the closest layer of gates to the top gate oxide in the three-layer gate design and screen electric fields to prevent unwanted regions of 2DEG/2DHG, effectively isolating the channels. The second layer includes the reservoir (blue) and plunger (red) gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confined potential. The chemical potential of each QD is primarily controlled by the corresponding plunger gate. The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs. By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
@@
BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays. The procedural flow of the single-button-press algorithm is depicted in Fig. 1(b).
BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments. Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes. Each of the six stages in BATIS involves at least one device measurement followed by data analysis to extract characteristics of a selected subset of gates, shown in Fig. 1(b) by gray rounded rectangles. The optional optical illumination stage is executed on an as-needed basis. Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state. Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [21].
The algorithm follows conditional logic visually represented with arrows of different colors in Fig. 1(b). The gray arrows indicate the procedural step in the sequence. Passing the test for the current stage, marked with yellow diamonds in Fig. 1(b), initiates the next stage, as indicated with green arrows. Failing a test can result in repeated measurements, device illumination, or termination of tuning, depending on the stage. The physical intuition behind each measurement, together with idealized and real-world data associated with each stage, are shown in Fig. 2.
@@
Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning. It also preconfigures the QD device for further use. The details of the flow of the algorithm are discussed in the following sections.
II.1 Initialization and First Leakage Test
The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections. The files also include details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage. These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to protect the device from having large damaging voltages applied. Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
For example, the functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table. The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments. This allows BATIS to remain device- and instrument-agnostic.
@@ II.1 Initialization and First Leakage Test
BATIS starts by checking the device for leakage to the fridge ground through the oxide or any possible means in the instrumentation setup. The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig. 1(b). The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig. 2(a). The desired result is a high resistance between all the connections.
Defects during the fabrication process can result in leakage. Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig. 2(a). Should leakage be detected, BATIS terminates and flags the device as unusable.
II.2 Global Accumulation
@@
For accumulation mode QD devices no 2DEG is present at zero applied gate voltage. Applying a non-zero bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage, see Fig. 1(b), is to determine the characteristic voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on.
Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal. Prior to fitting, the raw data is preprocessed to clean the signal. The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics followed by smoothing the data to remove noise from the signal. The processed signal is then fit to a function of the form
where is the center of the signal and describes the slope at . The final turn-on voltage is defined as
where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed.. In addition, the saturation of the curve is defined following Eq. (2), with . In cases where the saturation of the curve is not captured in the measurement, the last point along the sweep axis becomes . The first estimation of the maximum allowed voltage applied to each gate is defined as V.
@@
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 2(b), respectively. The turn-on process in real QD devices is not always straightforward, as even a good QD device might not turn on properly when first tested. If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig. 1(b).
@@
Illumination is initiated when the device fails to turn on or the turn-on voltage is much too high to be usable. After illumination, the global accumulation stage is re-initiated. If the preference is not to use illumination, BATIS allows skipping this step and going straight to repeating the global turn-on stage. Whether or not illumination should be used is an option that is pre-defined in the configuration file.
II.3 Accumulated Leakage Test

### v3

# Bootstrapping and autonomous tuning system for Si/SiGe multi-quantum dot devices

@@ Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. However, the increasing complexity of modern QD devices makes calibration and control—particularly at elevated temperatures—a bottleneck to progress, highlighting the need for robust and scalable autonomous solutions. A major hurdle arises from trapped charges within the oxide layers, which induce random offset voltage shifts on gate electrodes, with a standard deviation of approximately 83 of variation within state-of-the-art present-day devices. Efficient characterization and tuning of large arrays of QD qubits depend on choices of automated protocols. Here, we introduce a physically intuitive framework for a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing, formation of all current channels, and gate characterization in the presence of trapped charges. For forming the current channels, BATIS follows a non-standard approach that requires a single set of measurements regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring only minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, it will no longer be possible to tune up QD qubit arrays manually. The field of autotuning has arisen to address this challenge [9, 10].
@@
Modern semiconductor QD qubit devices have two critical features that enhance their performance while simultaneously making autotuning challenging and essential. First, the active regions of the device—the qubit QDs, charge sensors, and reservoirs—contain no doping. All charge carriers—electrons or holes—are induced by gate voltages and no current can flow until gate voltages are applied and carriers are accumulated. At the same time, the charges trapped within the gate-oxide heterostructure of the QD devices make it impossible to know a priori the precise turn-on voltages, i.e., the gate voltage threshold at which carrier accumulation begins. The device drift and hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) introduce additional variability in voltage configuration not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some host an array of QD qubits, while others provide a home for charge sensors. These channels are not separated until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve. However, when measuring a physical gate, a hysteresis can be observed.. The size of the QD introduces additional challenges: the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers. While spatially separated, the channels must remain coupled to ensure proper operation of proximal charge sensors on nearby QDs.
In this work, we introduce and demonstrate an intuitive bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure proper formation of all current channels. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including (critically) where current channels should be continuous and which gates should divide one channel from another. The output of BATIS is a comprehensive device characterization, specifying the gate voltage settings needed for QD chip operation, with clearly delineated current channels and identified pinch-off voltages.
@@
Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases [9]. For example, setting the device topology has been demonstrated for undoped accumulation-mode QD devices [12, 13]. Automated calibration of the charge state has been demonstrated for a modulation-doped QD device [14], an undoped SiMOS QD device [15], and undoped accumulation-mode QD devices [16].
@@
Methods for automatically detecting unintentional, spurious QDs have also been proposed [17]. More recently, virtual-gate control of the electrostatic potential landscape confining holes within a 2D QD array has been demonstrated [18].
Automated tuning of QD devices from a coldstart has been demonstrated for modulation-doped QD devices with predefined current channels [19, 20, 21] as well as for SiFinFETs and nanowires, which by design have a single current channel [21]. However, all these demonstrations assumed a prior knowledge of the device and a non-negligible device-specific preconfiguration. The first fully autonomous bootstrapping algorithm, intended for initiation on a pristine modulation-doped QD device, was introduced in Ref. [22]. However, the proposed algorithm is platform-specific and assumes the QD device to be functional. Autonomous characterization and bootstrapping for accumulation-mode QD devices have never been attempted.
BATIS addresses a key challenge in bootstrapping the QD devices: the diversity of device architectures and control layouts. All initialization tools proposed to date rely on rigid measurement sequences tailored to specific device designs, limiting their applicability across platforms. BATIS, in contrast, is designed to be platform-agnostic and thus can operate across a wide range of QD devices by dynamically adapting its procedures. While the procedural flow of BATIS is well-defined at a high level, the executable flow—the actual sequence of measurements and operations—is automatically adjusted in real time based on the device’s configuration and control wiring, as specified in a flexible configuration file. In other words, throughout execution, BATIS adjusts which steps to take and what outcomes to expect based on the device’s observed behavior. As tuning progresses, BATIS further tailors subsequent measurements to match the discovered characteristics of the system, ensuring efficient and device-specific calibration without manual intervention.
Importantly, the sample used to demonstrate BATIS is kept at around 1.3 , which is about 10 times warmer than other typical devices [23, 24]. High temperature operation is critical for practical quantum computer implementations [25]. However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [26]. The industry has demonstrated promising efforts in characterizing the performance of industry-manufactured spin qubit devices at 1.6 using the cryogenic 300- wafer prober [27, 28]. Autotuning of a single-electron transistor (SET) charge sensor in Si at 1.5 has also been demonstrated [29].
We demonstrate that BATIS performs reliably at temperatures above 1.0 and in the presence of trapped charges. It autonomously adapts to the specific QD layout and carries out critical initial tune-up stages, including identifying fabrication defects and characterizing both local and global current flow. These capabilities directly address emerging bottlenecks in scaling quantum semiconductor chips, making BATIS a powerful tool for enabling large-scale qubit implementations.
This paper is organized as follows: In Sec. II, we motivate the necessity of autotuning from the perspective of trapped charges. The bootstrap tuning procedure is described in detail in Sec. III. The experimental demonstration of BATIS to fully autonomously tune-up an accumulation mode Si/SixGe1-x QD device at 1.3 is described in Sec. IV. We conclude with a discussion of future directions in Sec. V.
II Trapped Charge: A Barrier to Robust Control
A primary challenge in automating current channel formation in QD devices stems from trapped charges within the QD array, which introduce substantial variability in the required gate voltages. Charge trapping refers to the immobilization of carriers at unknown locations due to, e.g., structural imperfections and defects on surfaces and interfaces of the heterostructure hosting the QDs. Stray charges can accumulate in any region of the heterostructure, including at buried interfaces, in the oxide layer at the top of the stack, and in particular, at the semiconductor-oxide interface, where there is an extremely high density of interface states [30].
Trapped charges in QD arrays present multiple challenges that degrade device performance [31]. First, they reduce the effective carrier concentration, which impedes transport and reduces mobility. In addition, trapped charges introduce temporal fluctuations in the local electric field, resulting in increased charge noise that compromises the stability and coherence of quantum states within the QDs. Over time, charge trapping can also lead to electrostatic disorder at the interface between the QD array and surrounding materials, further complicating device tunability and limiting the precision required for quantum control.
While trapped charges are largely locked into place as the device is cooled down from room temperature, they can nevertheless slowly drift over time, even at temperatures [32]. Moreover, these processes are not well controlled, leading to nonuniform charge accumulation that can vary from one cooldown to the next. As a result, significant and nonuniform voltage shifts may be required to offset the effects of trapped charges, and to achieve uniform gate operation across a device, even for nominally identical gate electrodes within a single device.
Such variability is undesirable for large-scale applications, and several techniques have been developed to mitigate this problem [33]. One such approach involves a simple “reset” of the trapped charges by optically illuminating the device in situ using a laser diode [34]. Illumination temporarily excites electrons above the bandgap of the host semiconductor, enabling the redistribution of immobile, unwanted trapped charges that otherwise remain fixed at low temperatures. Unfortunately, this procedure is not guaranteed to improve device performance and often needs to be repeated several times through trial and error, with different applied gate voltages.
@@ II Trapped Charge: A Barrier to Robust Control
The effect of trapped charges can be illustrated by simulating channel formation in a typical QD device. Figures 1(a-b) show a gate-defined, accumulation-mode, undoped QD device with multiple current channels, six designated regions for forming QDs (under the six plunger gates, shown in red), and reservoir gates (shown in blue) to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [35, 36].
@@
The screening gates (shown in yellow) form the bottommost layer in the three-layer gate architecture, positioned closest to the gate oxide. Their primary role is to screen electric fields, thereby preventing the unintentional formation of 2DEG or 2DHG regions and electrically isolating the intended transport channels. The second layer includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confinement potential. Each plunger gate primarily controls the chemical potential of the QD below it. The final layer consists of barrier gates (shown in gray), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs. By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
To model the effects of trapped charge, we introduce randomly varying charge densities on a tiled grid array at the oxide-semiconductor interface, with the density of each tile selected at random from a normal distribution with the standard deviation corresponding to a shift magnitude of , consistent with experimental observations for silicon QDs [28]. Each tile is , ensuring that all gates encompass multiple regions of randomly selected charge densities.
In the case where no random trapped charge is included, the ideal charge densities in the quantum well form coherent channels halfway between the screening gates, as indicated with the arrows in Fig. 1(b). In contrast, Figs. 1(c) and 1(e) show results obtained when randomly distributed trapped charges are included, resulting in current channels that flow beneath screening gates and pinched channels, respectively. Both problems can be resolved by careful adjustment of the various gate voltages to recover the coherent channels shown in Figs. 1(d) and 1(f), respectively.
While misdirected and pinched channels are easy to identify in simulations, they are much more difficult to diagnose in experiments. The issue of current flowing beneath the screening gate and unintentionally connecting multiple channels, as illustrated in Fig. 1(c), is intrinsic to 2D gate-defined systems and does not exist in 1D nanowire and nanotube devices [38, 39]. As such, it is of particular interest in this work. The recovery of misdirected current to form the desired coherent channels requires the simultaneous tuning of multiple gates. The protocols developed in this work are specially designed to address and characterize such 2D challenges.
III BATIS: Design and implementation
@@
BATIS is designed to automate the testing, characterization, and initialization of all gates and the forming of all necessary 1D current channels for linear QD arrays. The procedural flow of the single-button-press algorithm, depicted in Fig. 2, is rooted in the underlying device physics, making it both interpretable and robust, while also enabling adaptability across devices and scalability to arbitrary-length 1D arrays.
BATIS takes as input two configuration files describing the device design and certain generic specifications of the device and instruments. These files include detailed descriptions of all gates, the desired current channels, and the QD device connectivity. Additionally, they define a wire map of the experimental setup and a set of variables associated with each stage of the tuning process. These variables include instrumentation constraints, which prevent the algorithm from exploring unsupported operating ranges, and device safety limits, which protect the hardware from excessive or damaging voltages. Using this information, BATIS performs comprehensive diagnostics and characterization of each gate and automatically determines the appropriate configuration for the next stage of the tuning sequence, as depicted in Fig. 2.
The configuration files establish a link between abstract descriptions of device components and the operational language used within BATIS. For instance, the functionality of each gate is automatically inferred from the device architecture provided by the user, enabling BATIS to construct a gate lookup table. When combined with the wire map, this allows BATIS to automatically identify and perform the appropriate specialized measurements by linking the logical gate definitions to the physical measurement instruments. This abstraction layer ensures that BATIS remains agnostic to both the specific device layout and the instrumentation used.
The algorithm follows conditional logic, visually represented with boxes and arrows of different colors in Fig. 2. Each of the six stages in BATIS, indicated in Fig. 2 with gray rounded rectangles, involves at least one optimization of a device measurement or manipulation followed by data analysis to extract characteristics of a selected subset of gates. The measurements typical of each stage are detailed in Fig. 3, with the top row in each panel explaining the measurement, and the middle and bottom rows showing for each step a cartoon representation of the data (in each case, showing success in the left and failure on the right) and an example of measured experimental data, respectively.
The procedural steps in BATIS are indicated with gray arrows. Passing the functionality test for the current stage, marked with yellow diamonds in Fig. 2, initiates the next stage, as indicated with green arrows. Failing a test, indicated with red arrows, results in repeated measurements, device optical illumination to adjust the trapped charges, or bootstrapping termination, depending on the issues revealed by BATIS and the conclusions it draws based on the device physics.
@@
Once completed, BATIS produces a device diagnostic file that stores all information about each gate learned during tuning. When executed successfully, it verifies that all gates are functioning as intended and that isolated regions can be formed in well-defined 1D channels beneath the designated plunger gates. It also preconfigures the QD device for further use. In the following, we explain each state of the algorithm.
III.1 First Leakage Test
When initiated, BATIS first checks the device for leakage to the fridge ground through the oxide or any other possible means in the instrumentation setup. Leakage can be caused by, e.g., defects arising during the fabrication or packaging processes.
@@ III.1 First Leakage Test
The first leakage measurement happens before the electron accumulation, during the leakage test stage; see Fig. 2. The resistances of all the connections are measured with respect to each other and fridge ground, and they are compiled into a leakage matrix, as exemplified in the second and third row in Fig. 3(a). Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig. 3(a). The desired result is a high resistance between all the connections. Should leakage be detected, BATIS terminates and flags the device as unusable.
III.2 Global Accumulation
@@
For accumulation mode QD devices, no 2DEG is present at zero applied gate voltage. Applying a non-zero positive bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage is to determine the characteristic turn-on voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on; see Fig. 3(b). Since BATIS assumes no a priori knowledge about the device response, is determined through an iterative approach, with increased incrementally until BATIS is confident that is observed in the measured window.
@@
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 3(b), respectively. The turn-on process in real QD devices is not always straightforward, as even fully functional devices may fail to turn on correctly during initial testing. If the algorithm encounters a problem during global turn-on, it has a recovery pathway through optional illumination via a 780 light to adjust the microscopic configuration, magnitude, and sign of trapped charge, which can dramatically alter the device’s response to gate voltages [34]; see Fig. 2.
@@
After illumination, the global accumulation stage is re-initiated. If the preference is not to use illumination, BATIS allows skipping this step and going straight to repeating the global turn-on stage. Whether or not illumination should be used is an option that is pre-defined in the configuration file.
Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal. Prior to fitting, the raw data is preprocessed to clean the signal. The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics, followed by smoothing the data to remove noise from the signal and normalization. We then parameterize the response of the device using
where is the center of the step function and is the slope at . This function mimics the expected exponential behavior of the Fermi-Dirac distribution of electron energies with respect to gate voltages [40, 41, 42]. Similarly to the characteristics of the metal–oxide–semiconductor field-effect transistors (MOSFETs), an expected saturation will occur, leading to a roughly sigmoidal shape [43].
The final turn-on voltage is recorded as
where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed.. The safe maximum allowed voltage applied to each gate, , is then updated to better match the saturation of the curve, such that it is less than or equal to the original. This voltage is later used as the starting point for the characterization of the screening and reservoir pinch-offs.
III.3 Accumulated Leakage Test